### v1

# Directional anomaly detection

Abstract
Semi-supervised anomaly detection is based on the principle that potential anomalies are those records that look different from normal training data. However, in some cases we are specifically interested in anomalies that correspond to high attribute values (or low, but not both). We present two asymmetrical distance measures that take this directionality into account: ramp distance and signed distance. Through experiments on synthetic and real-life datasets we show that ramp distance performs as well or better than the absolute distance traditionally used in anomaly detection. While signed distance also performs well on synthetic data, it performs substantially poorer on real-life datasets. We argue that this reflects the fact that in practice, good scores on some attributes should not be allowed to compensate for bad scores on others.
Keywords:
a
nomaly detection, average localised proximity, distance measures, nearest neighbour distance, one-class classification, risk factors
1 Introduction
A defining characteristic of semi-supervised anomaly detection (also known as one-class classification [22]) is that the training set only contains normal data, but that we still want to create a model that can distinguish between normal and anomalous data. Potential reasons for working in this setting include not wanting to assume that the available examples of anomalies are representative of all anomalies, and wanting to flag all data that is different from normal data as a potential anomaly. Examples of application domains of anomality detection include failure detection in industrial settings and screening patients in healthcare.
In the present paper, we are specifically concerned with quantitative tabular data, where the training set consists of a representative sample of a certain normal target class. In this context, all semi-supervised anomaly detection algorithms are necessarily based on the principle that anomality increases with distance away from the normal training data in the feature space. However, there are circumstances in which we possess additional helpful domain knowledge. In this paper, we consider how best to deal with the knowledge that only relatively high values of an attribute should be indicative of anomality, not relatively low values (or vice-versa). For instance, certain attributes may correspond to higher strain of a machine, or encode known risk factors for patients. In such cases, we may want to decide that we are only interested in detecting anomalies that express high strain, or high risk factors, i.e., that we want to detect machine failure and at-risk patients, rather than underutilisation and exceptionally healthy patients.
To this end, we propose adaptations of two effective anomaly detection algorithms, Nearest Neighbour Distance (NND) and Average Localised Proximity (ALP), that take this directionality into account.
Other attempts to take domain knowledge into account include contextual anomaly detection, where anomality is conditioned on one or more contextual attributes [19, 14], and fair anomaly detection, where the goal is to ensure that anomalies follow the same distribution over one or more sensitive attributes as normal records [16, 26, 18]. However, to the best of our knowledge, this is the first work to investigate directional anomaly detection.
2 Background: NND and ALP
Semi-supervised anomaly detection algorithms, also known as one-class classifiers or data descriptors, take a training set $X$ consisting of normal data and learn a model that assigns an anomaly score to new instances. In the present paper, we will use two such algorithms: Nearest Neighbour Distance (NND) and Average Localised Proximity (ALP), both of which are directly based on distance measurements in the feature space.
2.1 NND
NND is one of the simplest anomaly detection algorithms and goes back to at least [10]. In NND, the anomaly score of a test record corresponds to the distance to its $k$th nearest neighbour in the training set. Despite its simplicity, NND performs surprisingly well across diverse datasets [12], making it an attractive baseline. In [11], a slight modification was proposed, taking the linearly weighted average of the first $k$ nearest neighbour distances. Thus, the anomality score of a test record $y$ becomes
$\sum_{i\leq k}w_{i}\cdot d_{i}(y),$ (1)
where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $w_{i}$ is the $i$th weight.
2.2 ALP
The idea of ALP [12] is to offset nearest neighbour distance against what is typical for normal data in that part of the feature space. It is based on the earlier form of localised nearest neighbour distance first proposed in [17] and [23].
ALP calculates a normality score in $[0,1]$ for a test record $y$ by taking the so-called weighted maximum of its localised proximity values:
$\operatorname*{\mathnormal{w}\,max\,}(\operatorname{lp}_{i}(y))_{i\leq k},$ (2)
for a choice of $k\geq 1$ and weight vector $w$, where the weighted maximum $\operatorname*{\mathnormal{w}\,max\,}$ is defined by
$\operatorname*{\mathnormal{w}\,max\,}X=\sum_{i\leq k}w_{i}\cdot X^{(i)},$ (3)
for any collection $X$, where $X^{(i)}$ is the $i$th largest element of $X$; and where $\operatorname{lp}_{i}(y)$ is the $i$th localised proximity of $y$ to the training set, defined by
$\operatorname{lp}_{i}(y)=\frac{D_{i}(y)}{D_{i}(y)+d_{i}(y)},$ (4)
where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $D_{i}(y)$ is the average $i$th nearest neighbour distance in the training set local to $y$, defined as
$D_{i}(y)=\sum_{j\leq l}w^{\prime}_{j}\cdot d_{i}(\operatorname{NN}_{j}(y)),$ (5)
for a choice of $l\geq 1$ and weights $w^{\prime}_{j}$, where $\operatorname{NN}_{j}(y)$ is the $j$th nearest neighbour of $y$ in the training set.
Note that $k$ fulfills a similar role as $k$ in weighted NND, whereas $l$ controls the amount of localisation (lower values correspond to more localisation).
2.3 Hyperparameters
For both NND and ALP, the values $k$ and $l$, the various weights and the distance measure are hyperparameters, which can in principle be tuned when there are anomalities available for validation purposes [13]. In general, however, we have to use sensible default values.
A typical choice for the distance $d$ is a form of the Minkowski $p$-distance:
$d(y,x)=\left(\sum_{j\leq m}\left\lvert y_{j}-x_{j}\right\rvert^{p}\right)^{\frac{1}{p}},$ (6)
for some value $p\geq 0$. The two most frequently used measures are $p=2$, which is Euclidean distance, and $p=1$, which is Boscovich distance, also known as city-block, Manhattan or rectilinear distance. For NND, Boscovich distance generally outperforms Euclidean distance [12, 11], while for ALP this is the standard choice anyway [12]. For both weighted NND and ALP, the default choice is to use linearly descending weights everywhere [12, 11]. For unweighted NND, $k=1$ is optimal [12], but for weighted NND slightly higher values like $k=8$ increase general performance [11]. For ALP, $k=5.5\log n$ and $l=6\log n$ have been established as optimal default values [12], where $n$ is the number of training records.
While not typically seen as a hyperparameter, both NND and ALP are also dependent on the relative scale of the attributes.
3 Directional anomaly detection
As discussed in the introduction, we want to find a way to take into account directional attributes, for which we should only interpret extreme values in one direction as anomalous. Without loss of generality, we will assume that high values correspond with anomality — any attributes for which the opposite applies can be transformed with a sign change. We will use the following potentially asymmetric distance measure: [^1] Note that while this introduces ramp distance and signed distance as variants of Boscovich distance, it is also possible to substitute ramp distance in the calculation of Minkowski distance (6) with other values of $p$, and in particular for Euclidean distance. For signed distance, adapting Minkowski distance requires more work.
$d(y,x)=\sum_{j\leq m}d_{j}(y_{i}-x_{i}),$ (7)
and consider three variants of the per-attribute distance $d_{j}(y_{i}-x_{i})$, listed in Table 1. Absolute distance is the baseline. It corresponds to the existing Boscovich distance, i.e. Minkowski distance (6) with $p=1$. Ramp distance is the result of applying the ramp function to the difference $y_{j}-x_{j}$. The ramp function is also known as the rectifier function, or rectified linear unit (ReLU), which is a popular activation function in neural networks [4]. Signed distance is simply the identity function applied to $y_{j}-x_{j}$.
Table 1: Variant approaches to directional attributes.
Variant $d_{j}(y_{i}-x_{i})$
Absolute $\left\lvert y_{j}-x_{j}\right\rvert$
Ramp $\max(0,y_{j}-x_{j})$
Signed $y_{j}-x_{j}$
All three variants agree for positive values of $y_{j}-x_{j}$; they interpret this as positive evidence for the anomality of $y$. They differ with respect to negative values of $y_{j}-x_{j}$, i.e., they differ in how to account for test values that are lower than the reference value in the training set. Absolute distance counts such low values as positive evidence for the anomality of $y$ just as it does high values, thereby ignoring the directionality of the attribute $j$. In contrast, signed distance interprets such low values as negative evidence for anomality, which can offset high values in other attributes. Finally, ramp distance stakes out the middle ground by interpreting such values as a lack of evidence either way.
Next, we consider how these variant distances affect NND and ALP. We can simply substitute the ramp distance into the overall distance for all directional attributes, and calculate the nearest neighbour distances according to the resulting measure.
For signed distance, the situation is not so straightforward. Note first that when all attributes are directional and we use signed distance, the nearest neighbour of any test record $y$ is the training instance $x$ that minimises $\sum_{i}(y_{i}-x_{i})$. This is the training instance with the largest attribute value sum $\sum_{i}x_{i}$, and this does not depend on the test instance. Analogously, all test instances share the same subsequently nearest neighbours. Consequently, we can disregard nearest neighbours altogether, and simply compare the attribute sums of test records. Note that we still use the training set to determine the right scale of each attribute. Nevertheless, because we no longer need to perform any nearest neighbour queries we end up with a significant computational simplification.
Figure 1: A dataset with one directional and one adirectional attribute, one test record $y$ and two training records $x$ and $x^{\prime}$. See text for discussion.
What should we do when we have a mixture of adirectional and directional attributes? One option is to add the signed distance from the directional attributes and the absolute distance from the adirectional attributes together to obtain a mixed distance between every test and training record and use this to identify the nearest neighbour distances of every test record. However, this would lead to the paradoxical situation illustrated by Fig. 1. The nearest neighbour of $y$ is $x^{\prime}$, since its mixed distance $4-6=-2$ is smaller than the mixed distance $2+0=2$ between $y$ and $x$. However, in terms of absolute distance, $y$ is clearly more similar to $x$, even if we only consider the adirectional attribute, and so it seems nonsensical to calculate the anomality of $y$ on the basis of its distance to $x^{\prime}$.
Therefore, we propose that for NND modified with signed distance, the right order of things is to seperately calculate the risk score for the directional attributes and the nearest neighbour distance for the adirectional attributes and add these together to obtain a single anomality score. Because of its effect on nearest neighbour calculations, we will not attempt to use signed distance with ALP.
Figure 2: Mean AUROC obtained with NND for synthetically generated datasets with Gaussian and Bernoulli attributes. $a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
Figure 3: Mean AUROC obtained with ALP for synthetically generated datasets with Gaussian and Bernoulli attributes. $a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
4 Experiments on synthetic datasets
In order to evaluate our proposal, we construct a series of synthetic datasets, each containing 1000 normal training records and 200 test records — 100 normal and 100 anomalous. Each record consists of 10 attributes with randomly generated values. We will consider two types of attributes. First, we generate datasets with continuous attributes, where both normal and anomalous values follow a Gaussian distribution with standard deviation 1, centred on 0 and $a$ respectively. We let $a$ vary between 0 and 1 in steps of 0.1. Secondly, we generate datasets with binary attributes, where normal and anomalous values follow Bernoulli distributions with $p=0.5-0.5\cdot b$ and $p=0.5+0.5\cdot b$ respectively, for which we let $b$ vary between 0 and 0.5 in steps of 0.05. For each value of $a$ and each value of $b$ we generate 100 datasets.
We evaluate anomaly detection performance using the area under the receiver operating characteristic (AUROC). This expresses the ability of an anomaly detector to separate anomalies from normal data. For this purpose, we monotonically transform the NND distance from (1) to an anomaly score in $[0,1]$ with:
$a\longmapsto\frac{1}{2}\cdot\frac{a}{\left\lvert a\right\rvert+1}+\frac{1}{2}.$ (8)
We do this for absolute distance, ramp distance and signed distance. The results for NND are displayed in Figure 2. For the Gaussian attributes, signed and ramp distance perform substantially better than absolute distance, for small and large $k$, in line with our directional hypothesis. Signed distance performs slightly better than ramp distance, but the difference is small. For Bernoulli attributes, the picture looks a bit different. At $k=1$, signed distance outperforms absolute distance, but ramp distance performs worse than absolute distance. At $k=8$, the gap between signed and absolute distance has tightened considerably, and ramp distance now sits in between. At $k=100$, signed and ramp distance have become indistinguishable in terms of performance, and absolute distance is only very slightly behind.
For ALP, the results for the default hyperparameter values $k=5.5\log 1000=38$ and $l=6\log 1000=41$ are displayed in Table 3. For both Gaussian and Bernoulli attributes, ramp distance outperforms absolute distance, although more clearly so for Gaussian attributes.
Taken together, these results suggest that when we know for a given attribute that anomalies have a higher mean value than normal data, we should indeed use signed or ramp distance rather than absolute distance. Between these two variants, signed distance appears to have the advantage with NND, but we will see in the next section that the situation is different for real-life datasets.
Table 2: Real-life datasets used in the experiment. $m$: number of attributes.
Dataset Source Records Anomalies $m$
ai4i2020 [15] 10 000 Simulated machine operation records 339 Five different failure modes (not distinguished by the label) 6
diabetes-risk [8] 520 Patients with diabetes symptoms 320 Patients actually diagnosed with diabetes 14
fertility [5] 100 Sperm samples 12 ‘Altered’ samples 8
heart-failure [1] 299 Heart patients 96 Patients who died in the follow-up period 11
phishing-websites [24] 11 055 Websites 4898 Phishing websites 30
post-operative [2] 87 Patients who have undergone surgery 63 Patients who had to stay in hospital 8
qualitative-bankruptcy [9] 250 Companies 107 Companies that went bankrupt 6
south-german-credit [6] 1000 Credits provided by a bank 300 Bad credits 20
thoraric-surgery [27] 470 Primary lung cancer patients who underwent major lung resections 70 Patients who died within one year 16
wdbc [21] 569 Images of aspirated breast tumor samples 212 Malignant tumors 30
wisconsin [25] 683 Fine-needle aspirates of breast tumors 239 Malignant tumors 9
wpbc [20] 138 Patients who have undergone surgery for an invasive malignant breast cancer tumor with no evidence of distant metastases 28 Patients who experienced recurrence within two years 32
Table 3: Mean cross-validation AUROC for the absolute, ramp and signed distance variants. Bold: highest value (before rounding) for NND and ALP, respectively.
Dataset NND ALP
Absolute Ramp Signed Absolute Ramp
ai4i2020 0.823 0.922 0.724 0.877 0.924
diabetes-risk 0.971 0.923 0.716 0.895 0.926
fertility 0.602 0.653 0.540 0.581 0.636
heart-failure 0.715 0.769 0.735 0.734 0.766
phishing-websites 0.901 0.927 0.804 0.927 0.936
post-operative 0.476 0.504 0.557 0.459 0.484
qualitative-bankruptcy 1.000 1.000 0.998 1.000 1.000
south-german-credit 0.648 0.718 0.683 0.648 0.714
thoraric-surgery 0.597 0.624 0.583 0.634 0.621
wdbc 0.950 0.976 0.969 0.957 0.981
wisconsin 0.995 0.994 0.995 0.872 0.995
wpbc 0.570 0.625 0.633 0.537 0.654
5 Experiments on real-life datasets
We will now evaluate our proposal on a number of real-life datasets from the UCI repository [3] that can be approached as directional anomaly detection problems. The main properties of these datasets are listed in Table 2. The attributes of these datasets express machine operating conditions (ai4i2020), medical symptoms, comorbidities and lifestyle factors (diabetes-risk, fertility, heart-failure, post-operative and thoraric-surgery), tumor characteristics (wdbc, wisconsin and wpbc), and risk indicators provided by experts (phishing-websites, qualitative-bankrupty and south-german-credit). ai4i2020 and post-operative contain a mixture of directional and adirectional attributes, the other datasets only have directional attributes.
As in the previous section, we use AUROC to evaluate anomaly detection performance. We perform 5-fold cross-validation on the normal records from each dataset, creating at each iteration a test set by combining one fifth of the normal records with all anomalous records. For NND, we use linear weights and fix $k=8$. We rescale all attributes by subtracting the midhinge and dividing by the semi-interquartile range of the normal training values, such that the interquartile range becomes $[-1,1]$ in the training data.
Table 3 lists the mean cross-validation AUROC obtained by the various variants for each dataset. The first thing to note is that for NND, unlike what we saw for synthetic data in the previous section, ramp distance generally performs (much) better than signed distance. The only datasets for which signed distance performs better are post-operative and wpbc, for which the AUROC scores are very low anyway. When we perform a one-side Wilcoxon signed-rank test, we obtain $p=0.021$. We can conclude from this that real-life anomaly detection problems are different from our synthetic datasets in one important aspect: an unexpectedly low score on one risk factor does not compensate for high scores on other risk factors.
However, these experiments do confirm the usefulness of directional anomaly detection, because ramp distance performs about as well or better than absolute distance. For NND, the signed-rank test gives $p=0.011$, or $p=0.023$ if we apply Holm-Bonferroni correction [7] for multiple testing, while for ALP it gives $p=0.0029$.
There is one notable exception: diabetes-risk for NND. A potential explanation for this is the fact that the attributes of this dataset encode diabetes symptoms, and that moreover, the normal records do not represent average healthy people, but non-diabetes patients who nevertheless display symptoms of diabetes. Consequently, this may undermine the directionality of this dataset — diabetes patients may not necessarily have more symptoms than this particular group of other patients, but rather different combinations of symptoms, which would make absolute distance a better fit than ramp distance.
Indeed, when we look at the mean attribute values of normal and anomalous records in this dataset, we find that for five of the fourteen attributes, the anomalous records (diabetes patients) have lower or only slightly higher mean values than the normal records. If we rerun the experiment while treating these five attributes as non-directional, the AUROC obtained by ramp distance matches that of absolute distance.
6 Conclusion
In this paper, we have introduced directional anomaly detection, a new problem setting wherein only high (or only low) values of certain attributes are indicative of anomality, and which can therefore be viewed as risk factors. We have proposed two different ways of adapting distance-based anomaly detection algorithms to make use of this knowledge: ramp distance and signed distance. The difference between these two variants corresponds to a choice in how a practitioner wants to interpret an unexpectedly low value on a given risk factor. With ramp distance, such low values are simply discounted, whereas with signed distance, they contribute negative evidence against anomality. Both variants contrast with absolute distance, the non-directional baseline wherein such low values contribute positive evidence for anomality.
In an experiment with synthetically generated Gaussian and Bernoulli data, we found that signed distance is able to achieve slightly higher anomaly detection than ramp distance. However, in a subsequent experiment with real-life datasets, ramp distance performed substantially better. We conclude from this that in typical real-life use cases, unexpectedly low values on risk factors should not compensate for high values on other risk factors.
We also found that ramp distance generally performed as good or better than absolute distance. Therefore, we recommend the use of ramp distance when a practitioner knows that some of the attributes in their dataset are really risk factors.
Finally, we will end by briefly considering what it means if, for a given dataset, ramp distance should nevertheless result in worse anomaly detection performance than absolute distance. One possibility is that the directional hypothesis is false for one or more of the attributes, that there is a good reason after all why (some) anomalies should have lower values than normal. This is in fact what we saw with one of the real-life datasets that we evaluated, the only dataset where absolute distance clearly performed better than ramp distance for NND.
The other possibility is that there is no clear explanation, no plausible causal relation linking a lower attribute value to a higher risk of abnormality. In that case using ramp distance may still be preferable, because it better matches the domain knowledge of the practitioner, resulting in a more interpretable prediction model.
Acknowledgements
This publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO).

### v2

@@
-Directional anomaly detection
+Monotonic anomaly detection
 Abstract
@@
 However, in some cases we are specifically interested in anomalies that correspond to high attribute values (or low, but not both).
-We present two asymmetrical distance measures that take this directionality into account: ramp distance and signed distance.
-Through experiments on synthetic and real-life datasets we show that ramp distance performs as well or better than the absolute distance traditionally used in anomaly detection.
+We present two asymmetrical distance measures that take this monotonicity into account: ramp distance and signed distance.
+Through experiments on synthetic and real-life datasets, we show that ramp distance increases anomaly detection performance over the traditional absolute distance.
 While signed distance also performs well on synthetic data, it performs substantially poorer on real-life datasets.
-We argue that this reflects the fact that in practice, good scores on some attributes should not be allowed to compensate for bad scores on others.
+We argue that this is a consequence of the fact that when using signed distance, low values of certain attributes automatically compensate for high values of other attributes, such that anomaly detection is reduced to counting the total attribute value sum, which is too simplistic in practice.
 Keywords:
 a
-nomaly detection, average localised proximity, distance measures, nearest neighbour distance, one-class classification, risk factors
+nomaly detection, distance measures, monotonicity, nearest neighbours, one-class classification, risk factors
 1 Introduction
-A defining characteristic of semi-supervised anomaly detection (also known as one-class classification [22]) is that the training set only contains normal data, but that we still want to create a model that can distinguish between normal and anomalous data.
-Potential reasons for working in this setting include not wanting to assume that the available examples of anomalies are representative of all anomalies, and wanting to flag all data that is different from normal data as a potential anomaly.
+A defining characteristic of semi-supervised anomaly detection (also known as one-class classification [22]) is that the training set only contains normal records, but that we still want to create a model that can distinguish between normal records and anomalies.
+Potential reasons for working in this setting include not wanting to assume that the available examples of anomalies are representative of all anomalies, and wanting to flag all new records that deviate from normal data as potential anomalies.
 Examples of application domains of anomality detection include failure detection in industrial settings and screening patients in healthcare.
@@
 For instance, certain attributes may correspond to higher strain of a machine, or encode known risk factors for patients.
-In such cases, we may want to decide that we are only interested in detecting anomalies that express high strain, or high risk factors, i.e., that we want to detect machine failure and at-risk patients, rather than underutilisation and exceptionally healthy patients.
-To this end, we propose adaptations of two effective anomaly detection algorithms, Nearest Neighbour Distance (NND) and Average Localised Proximity (ALP), that take this directionality into account.
-Other attempts to take domain knowledge into account include contextual anomaly detection, where anomality is conditioned on one or more contextual attributes [19, 14], and fair anomaly detection, where the goal is to ensure that anomalies follow the same distribution over one or more sensitive attributes as normal records [16, 26, 18].
-However, to the best of our knowledge, this is the first work to investigate directional anomaly detection.
-2 Background: NND and ALP
+In such cases, we may only be interested in detecting anomalies that express high strain, or high risk factors, i.e., machine failure and at-risk patients, rather than underutilisation and exceptionally healthy patients.
+We call this setting monotonic anomaly detection, because our domain knowledge places a monotonicity constraint on certain attributes and the prediction.
+Alternatively, we may think of this as directional anomaly detection, because the goal is to detect anomalies that lie in a certain direction from the normal data in the attribute space.
+When applied to such a monotonic anomaly detection problem, an ordinary anomaly detection algorithm will needlessly flag records that represent underutilisation or exceptionally healthy patients.
+But the real problem is more fundamental.
+Existing anomaly detection algorithms will flag any record that deviates sufficiently from the normal records, due to a combination of relatively high values for some attributes and relatively low values for others.
+This is problematic if we believe that relatively low values should not be indicative of anomality.
+Consequently, we propose that a correct use of monotonic domain knowledge requires that we adapt existing algorithms by modifying the contribution of individual attribute values to the anomaly score.
+Specifically, we have to ask what the contribution of a relatively low attribute value of a test record should be.
+Logically, there are three possibilities: relatively low attribute values can be interpreted as positive evidence of anomality, as negative evidence against anomality, and as neither positive nor negative evidence.
+The first of these three alternatives corresponds to ordinary anomality detection, because it makes no difference between relatively low and relatively high attribute values — both are treated as positive evidence of anomality.
+In contrast, the second and third option represent two alternative monotonic approaches, because they treat low and high attribute values differently.
+In the present paper, we investigate how distance-based algorithms can be modified for monotonic anomaly detection.
+Specifically, we will consider Centre Distance (CD), Nearest Neighbour Distance (NND), Local Outlier Factor (LOF) and Average Localised Proximity (ALP).
+The reason for focusing on these algorithms is twofold.
+Firstly, they are all based on the distance between records, and we can make them monotonic by choosing an appropriate asymmetric distance measure.
+Secondly, these are some of the best-performing algorithms in the literature.
+In an experimental evaluation across 246 semi-supervised anomaly detection problems [12], ALP, NND and LOF placed first, third and fourth.
+In Section 2, we briefly mention some related work, and in Section 3 we describe the CD, NND, LOF and ALP algorithms.
+Next, we present our proposed asymmetric distance measures that enable directional anomaly detection in Section 4.
+Finally, we describe our experimental setup in Section 5 and present the results in Section 6, before concluding in Section 7.
+We also include Appendix 0.A with an analysis of the hyperparameter choice $k$ for NND.
+2 Related work
+There exists a substantial literature on monotonic classification [3].
+However, because approaches proposed in that context rely on a representative sample from each class in the training set, they cannot be applied to semi-supervised anomaly detection.
+The closest related work to the present paper is the proposal on direction-related anomaly detection [25].
+The problem setting in that work is a form of ordinal regression, where anomalies are the records with the highest response value, which are not represented in the training set.
+The proposed approach is to use an ordinal regression algorithm to learn a decision boundary beyond which lie the anomalies.
+Thus, the directionality in that setting is determined by the response values, rather than the independent attributes.
+Other attempts to incorporate domain knowledge in anomaly detection include contextual anomaly detection, where anomality is conditioned on one or more contextual attributes [19, 14], and fair anomaly detection, where the goal is to ensure that anomalies follow the same distribution over one or more sensitive attributes as normal records [16, 27, 18].
+3 Anomaly detection algorithms
 Semi-supervised anomaly detection algorithms, also known as one-class classifiers or data descriptors, take a training set $X$ consisting of normal data and learn a model that assigns an anomaly score to new instances.
-In the present paper, we will use two such algorithms: Nearest Neighbour Distance (NND) and Average Localised Proximity (ALP), both of which are directly based on distance measurements in the feature space.
+In the present paper, we will consider four such algorithms: Centre Distance (CD), Nearest Neighbour Distance (NND), Local Outlier Factor (LOF) and Average Localised Proximity (ALP).
+3.1 CD
+CD is arguably the simplest possible anomaly detection algorithm.
+It requires choosing a centre $c$, a single value such as the marginal mean or median that is representative of the whole training set.
+The anomaly score of a test record $y$ is its distance to $c$.
+Due to its simplicity, CD is a useful baseline.
 3.2 NND
-NND is one of the simplest anomaly detection algorithms and goes back to at least [10].
-In NND, the anomaly score of a test record corresponds to the distance to its $k$th nearest neighbour in the training set.
-Despite its simplicity, NND performs surprisingly well across diverse datasets [12], making it an attractive baseline.
+NND is also very simple, and goes back to at least [10].
+The anomaly score of a test record corresponds to the distance to its $k$th nearest neighbour in the training set.
+Despite its simplicity, NND performs surprisingly well across diverse datasets [12], making it an attractive default choice for anomaly detection.
 In [11], a slight modification was proposed, taking the linearly weighted average of the first $k$ nearest neighbour distances.
-Thus, the anomality score of a test record $y$ becomes
+Thus, the anomality score of a test record $y$ becomes:
 $\sum_{i\leq k}w_{i}\cdot d_{i}(y),$ (1)
 where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $w_{i}$ is the $i$th weight.
+3.3 LOF
+The idea behind LOF is to estimate the density of the normal data at the location of a test record $y$, and to compare this to the density at the location of its nearest neighbours.
+The anomality score of $y$ corresponds inversely to this relative density.
+Concretely, it is calculated as follows:
+$\frac{1}{k}\sum_{j\leq k}\frac{\operatorname{lrd}_{k}(\operatorname{NN}_{j}(y))}{\operatorname{lrd}_{k}(y)},$ (2)
+for a choice of $k\geq 1$, and a distance measure $d$, where $\operatorname{NN}_{j}(y)$ is the $j$th nearest neighbour of $y$ in the training set, and $\operatorname{lrd}_{k}$ is the so-called local reachability density, which in turn corresponds inversely to the distance between a record and its neighbours:
+$\operatorname{lrd}_{k}(y)=\frac{1}{\frac{1}{k}\sum_{i\leq k}\operatorname{rd}_{k}(y,\operatorname{NN}_{i}(y))},$ (3)
+where $\operatorname{rd}_{k}$ is the so-called reachability distance:
+$\operatorname{rd}_{k}(y,x)=\max(d(y,x),d_{k}(x)),$ (4)
+where $d_{k}(x)$ is the $k$th nearest neighbour distance of $x$ in the training set.
+The motivation for using the asymmetric reachability distance rather than ordinary distance is that this results in a smoother density function.
+LOF is a well-known and widely-used algorithm. Nevertheless, when evaluated across a wide range of real-life datasets, it failed to beat the simpler NND algorithm [12].
 3.4 ALP
@@
 $\operatorname*{\mathnormal{w}\,max\,}(\operatorname{lp}_{i}(y))_{i\leq k},$ (5)
-for a choice of $k\geq 1$ and weight vector $w$, where the weighted maximum $\operatorname*{\mathnormal{w}\,max\,}$ is defined by
+for a choice of $k\geq 1$ and weight vector $w$, where the weighted maximum $\operatorname*{\mathnormal{w}\,max\,}$ is defined by:
 $\operatorname*{\mathnormal{w}\,max\,}X=\sum_{i\leq k}w_{i}\cdot X^{(i)},$ (6)
-for any collection $X$, where $X^{(i)}$ is the $i$th largest element of $X$; and where $\operatorname{lp}_{i}(y)$ is the $i$th localised proximity of $y$ to the training set, defined by
+for any collection $X$, where $X^{(i)}$ is the $i$th largest element of $X$; and where $\operatorname{lp}_{i}(y)$ is the $i$th localised proximity of $y$ to the training set, defined by:
 $\operatorname{lp}_{i}(y)=\frac{D_{i}(y)}{D_{i}(y)+d_{i}(y)},$ (7)
-where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $D_{i}(y)$ is the average $i$th nearest neighbour distance in the training set local to $y$, defined as
+where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $D_{i}(y)$ is the average $i$th nearest neighbour distance in the training set local to $y$, defined as:
 $D_{i}(y)=\sum_{j\leq l}w^{\prime}_{j}\cdot d_{i}(\operatorname{NN}_{j}(y)),$ (8)
@@
 Note that $k$ fulfills a similar role as $k$ in weighted NND, whereas $l$ controls the amount of localisation (lower values correspond to more localisation).
-2.3 Hyperparameters
-For both NND and ALP, the values $k$ and $l$, the various weights and the distance measure are hyperparameters, which can in principle be tuned when there are anomalities available for validation purposes [13].
-In general, however, we have to use sensible default values.
-A typical choice for the distance $d$ is a form of the Minkowski $p$-distance:
-$d(y,x)=\left(\sum_{j\leq m}\left\lvert y_{j}-x_{j}\right\rvert^{p}\right)^{\frac{1}{p}},$ (6)
-for some value $p\geq 0$.
-The two most frequently used measures are $p=2$, which is Euclidean distance, and $p=1$, which is Boscovich distance, also known as city-block, Manhattan or rectilinear distance.
-For NND, Boscovich distance generally outperforms Euclidean distance [12, 11], while for ALP this is the standard choice anyway [12].
-For both weighted NND and ALP, the default choice is to use linearly descending weights everywhere [12, 11].
-For unweighted NND, $k=1$ is optimal [12], but for weighted NND slightly higher values like $k=8$ increase general performance [11].
-For ALP, $k=5.5\log n$ and $l=6\log n$ have been established as optimal default values [12], where $n$ is the number of training records.
-While not typically seen as a hyperparameter, both NND and ALP are also dependent on the relative scale of the attributes.
-3 Directional anomaly detection
-As discussed in the introduction, we want to find a way to take into account directional attributes, for which we should only interpret extreme values in one direction as anomalous.
-Without loss of generality, we will assume that high values correspond with anomality — any attributes for which the opposite applies can be transformed with a sign change.
-We will use the following potentially asymmetric distance measure: [^1] Note that while this introduces ramp distance and signed distance as variants of Boscovich distance, it is also possible to substitute ramp distance in the calculation of Minkowski distance (6) with other values of $p$, and in particular for Euclidean distance.
-For signed distance, adapting Minkowski distance requires more work.
-$d(y,x)=\sum_{j\leq m}d_{j}(y_{i}-x_{i}),$ (7)
-and consider three variants of the per-attribute distance $d_{j}(y_{i}-x_{i})$, listed in Table 1.
-Absolute distance is the baseline.
-It corresponds to the existing Boscovich distance, i.e.
-Minkowski distance (6) with $p=1$.
-Ramp distance is the result of applying the ramp function to the difference $y_{j}-x_{j}$.
-The ramp function is also known as the rectifier function, or rectified linear unit (ReLU), which is a popular activation function in neural networks [4].
-Signed distance is simply the identity function applied to $y_{j}-x_{j}$.
-Table 1: Variant approaches to directional attributes.
-Variant $d_{j}(y_{i}-x_{i})$
-Absolute $\left\lvert y_{j}-x_{j}\right\rvert$
-Ramp $\max(0,y_{j}-x_{j})$
-Signed $y_{j}-x_{j}$
-All three variants agree for positive values of $y_{j}-x_{j}$; they interpret this as positive evidence for the anomality of $y$.
-They differ with respect to negative values of $y_{j}-x_{j}$, i.e., they differ in how to account for test values that are lower than the reference value in the training set.
-Absolute distance counts such low values as positive evidence for the anomality of $y$ just as it does high values, thereby ignoring the directionality of the attribute $j$.
-In contrast, signed distance interprets such low values as negative evidence for anomality, which can offset high values in other attributes.
-Finally, ramp distance stakes out the middle ground by interpreting such values as a lack of evidence either way.
-Next, we consider how these variant distances affect NND and ALP.
-We can simply substitute the ramp distance into the overall distance for all directional attributes, and calculate the nearest neighbour distances according to the resulting measure.
-For signed distance, the situation is not so straightforward.
-Note first that when all attributes are directional and we use signed distance, the nearest neighbour of any test record $y$ is the training instance $x$ that minimises $\sum_{i}(y_{i}-x_{i})$.
-This is the training instance with the largest attribute value sum $\sum_{i}x_{i}$, and this does not depend on the test instance.
-Analogously, all test instances share the same subsequently nearest neighbours.
-Consequently, we can disregard nearest neighbours altogether, and simply compare the attribute sums of test records.
-Note that we still use the training set to determine the right scale of each attribute.
-Nevertheless, because we no longer need to perform any nearest neighbour queries we end up with a significant computational simplification.
-Figure 1: A dataset with one directional and one adirectional attribute, one test record $y$ and two training records $x$ and $x^{\prime}$. See text for discussion.
-What should we do when we have a mixture of adirectional and directional attributes?
-One option is to add the signed distance from the directional attributes and the absolute distance from the adirectional attributes together to obtain a mixed distance between every test and training record and use this to identify the nearest neighbour distances of every test record.
-However, this would lead to the paradoxical situation illustrated by Fig.
-1.
-The nearest neighbour of $y$ is $x^{\prime}$, since its mixed distance $4-6=-2$ is smaller than the mixed distance $2+0=2$ between $y$ and $x$.
-However, in terms of absolute distance, $y$ is clearly more similar to $x$, even if we only consider the adirectional attribute, and so it seems nonsensical to calculate the anomality of $y$ on the basis of its distance to $x^{\prime}$.
-Therefore, we propose that for NND modified with signed distance, the right order of things is to seperately calculate the risk score for the directional attributes and the nearest neighbour distance for the adirectional attributes and add these together to obtain a single anomality score.
-Because of its effect on nearest neighbour calculations, we will not attempt to use signed distance with ALP.
-Figure 2: Mean AUROC obtained with NND for synthetically generated datasets with Gaussian and Bernoulli attributes.
-$a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
-Figure 3: Mean AUROC obtained with ALP for synthetically generated datasets with Gaussian and Bernoulli attributes.
-$a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
-4 Experiments on synthetic datasets
-In order to evaluate our proposal, we construct a series of synthetic datasets, each containing 1000 normal training records and 200 test records — 100 normal and 100 anomalous.
+In the large experiment on real-life datasets [12], ALP came out on top.
+4 Monotonic anomaly detection
+As discussed in the introduction, we want to find a way to take into account monotonic attributes, for which we should only interpret extreme values in one direction as anomalous.
+Without loss of generality, we will assume that high values correspond to anomality — any attributes for which the opposite applies can be transformed with a sign change.
+4.1 Asymmetric distance
+The anomaly detection algorithms described in Section 3 are all based on a choice of distance measure $d$ between various records.
+For tabular data, a typical choice for the distance $d$ is a form of Minkowski $p$-distance:
+$d(y,x)=\Big{(}\sum_{i\leq m}\lvert y_{i}-x_{i}\rvert^{p}\Big{)}^{\frac{1}{p}},$ (9)
+for some $p\in(0,\infty)$.
+[^1] In addition, Minkowski $\infty$-distance, also known as Chebyshev distance, is the limit case $p\rightarrow\infty$, which resolves to $d(y,x)=\max_{i\leq m}\left\lvert y_{i}-x_{i}\right\rvert$.
+In particular, the two most frequently used choices are 1-distance and 2-distance.2221-distance is also known as Boscovich, city-block or Manhattan distance.
+2-distance is also known as Euclidean distance.
+Minkowski $p$-distance is a function of the per-attribute differences $y_{i}-x_{i}$.
+In the anomaly detection algorithms described in Section 3, the distance is applied to test records $y$ and training records $x$.
+When $y_{i}$ is larger than $x_{i}$, this contributes positive evidence to the anomality of $y$.
+And because the Minkowski $p$-distance takes the absolute value of the per-attribute differences, values of $y_{i}$ that are smaller than $x_{i}$ likewise contribute positive evidence for the anomality of $y$.
+However, when we are dealing with a monotonic attribute, small values of $y_{i}$ are not an indication of anomality.
+Logically, there are two alternative ways that values of $y_{i}$ smaller than $x_{i}$ can be interpreted.
+By not taking the absolute value, but simply using the signed difference $y_{i}-x_{i}$, we interpret small values of $y_{i}$ as negative evidence against anomality.
+And by applying the ramp function [^3] This is also known as the rectifier function, or rectified linear unit (ReLU), and is a popular activation function in neural networks [5].
+$\max(0,y_{i}-x_{i})$, we do not interpret small values of $y_{i}$ as providing evidence either for or against anomality.
+In general, a dataset may contain a mixture of monotonic and ordinary attributes. Therefore, we propose the following generalised, potentially asymmetric form of Minkowski $p$-distance:
+[…]
+[…]
+We will use absolute distance for all ordinary attributes, and compare the use of absolute, ramp and signed distance for monotonic attributes.
+Using absolute distance for monotonic attributes is the baseline, since this corresponds to traditional anomaly detection.
+Table 1: Variant approaches to monotonic attributes.
+Variant $\delta_{i}(y_{i}-x_{i})$ Interpretation of negative values
+Absolute $\left\lvert y_{i}-x_{i}\right\rvert$ Evidence for anomality
+Ramp $\max(0,y_{i}-x_{i})$ No evidence either way
+Signed $y_{i}-x_{i}$ Evidence against anomality
+4.2 Anomaly detection with signed distance
+Whereas ramp distance can be straightforwardly substituted into the definitions of the anomaly detection algorithms described in Section 3, the use of signed distance leads to complications.
+To simplify the analysis, we will use 1-distance, and consider a dataset that consists entirely of monotonic attributes, which is the case for the majority of datasets in our experiments.
+For CD, we find that when we use signed distance, the choice of centre $c\in\mathbb{R}^{m}$ does not actually matter.
+By definition, the anomaly score of a test record $y\in\mathbb{R}^{m}$ is $\sum_{i\leq m}(y_{i}-c_{i})$.
+Therefore, a change in centre from $c$ to $c^{\prime}$ simply induces a shift of all anomaly scores by a constant $\sum_{i\leq m}(c_{i}-c^{\prime}_{i})$.
+Effectively, the anomaly score of $y$ is simply the sum of its attribute values, i.e.
+the total risk of a record.
+[^4] This reduces the role of the training set to scaling, i.e.
+to determine the relative weight of each risk factor.
+Furthermore, the nearest neighbour of $y$ is by definition the training record $x$ that minimises $\sum_{i\leq m}(y_{i}-x_{i})$, which is the training record $x$ with the highest attribute value sum $\sum_{i\leq m}x_{i}$.
+This is independent of $y$, meaning that we lose the local character of nearest neighbour algorithms.
+In particular, NND with signed distance is equivalent to CD, since we saw above that the choice of centre does not fundamentally matter for the anomaly score of CD, and we can choose $x$ to be this centre.
+In light of this, we can only evaluate signed distance for CD.
+5 Experimental setup
+We will conduct a series of experiments on synthetic and real-life datasets from the UCI repository [4] to evaluate whether the approach proposed in this paper leads to better anomaly detection performance on datasets with monotonic attributes.
+Concretely, we will compare the use of absolute distance (ordinary anomaly detection) with ramp distance (monotonic anomaly detection), as well as signed distance for the CD algorithm.
+5.1 Evaluation
+We evaluate anomaly detection performance using the area under the receiver operating characteristic (AUROC).
+This expresses the ability of an anomaly detector to separate anomalies from normal data.
+For this purpose, we monotonically transform the NND score from (1) and the LOF score (2) to an anomaly score in $[0,1]$ with:
+$a\longmapsto\frac{a}{\left\lvert a\right\rvert+1}.$ (11)
+Because the CD score can be negative when using signed distance, we instead use the following function to transform it to an anomaly score in $[0,1]$:
+$a\longmapsto\frac{1}{2}\cdot\frac{a}{\left\lvert a\right\rvert+1}+\frac{1}{2}.$ (12)
+5.2 Synthetic datasets
+The synthetic datasets contain 1000 normal training records and 200 test records — 100 normal and 100 anomalous.
 Each record consists of 10 attributes with randomly generated values.
@@
 We let $a$ vary between 0 and 1 in steps of 0.1.
-Secondly, we generate datasets with binary attributes, where normal and anomalous values follow Bernoulli distributions with $p=0.5-0.5\cdot b$ and $p=0.5+0.5\cdot b$ respectively, for which we let $b$ vary between 0 and 0.5 in steps of 0.05.
-For each value of $a$ and each value of $b$ we generate 100 datasets.
-We evaluate anomaly detection performance using the area under the receiver operating characteristic (AUROC).
-This expresses the ability of an anomaly detector to separate anomalies from normal data.
-For this purpose, we monotonically transform the NND distance from (1) to an anomaly score in $[0,1]$ with:
-$a\longmapsto\frac{1}{2}\cdot\frac{a}{\left\lvert a\right\rvert+1}+\frac{1}{2}.$ (8)
-We do this for absolute distance, ramp distance and signed distance.
-The results for NND are displayed in Figure 2.
-For the Gaussian attributes, signed and ramp distance perform substantially better than absolute distance, for small and large $k$, in line with our directional hypothesis.
-Signed distance performs slightly better than ramp distance, but the difference is small.
-For Bernoulli attributes, the picture looks a bit different.
-At $k=1$, signed distance outperforms absolute distance, but ramp distance performs worse than absolute distance.
-At $k=8$, the gap between signed and absolute distance has tightened considerably, and ramp distance now sits in between.
-At $k=100$, signed and ramp distance have become indistinguishable in terms of performance, and absolute distance is only very slightly behind.
-For ALP, the results for the default hyperparameter values $k=5.5\log 1000=38$ and $l=6\log 1000=41$ are displayed in Table 3.
-For both Gaussian and Bernoulli attributes, ramp distance outperforms absolute distance, although more clearly so for Gaussian attributes.
-Taken together, these results suggest that when we know for a given attribute that anomalies have a higher mean value than normal data, we should indeed use signed or ramp distance rather than absolute distance.
-Between these two variants, signed distance appears to have the advantage with NND, but we will see in the next section that the situation is different for real-life datasets.
+Second, we generate datasets with binary attributes, where normal and anomalous values follow Bernoulli distributions with $p=0.5-0.5\cdot b$ and $p=0.5+0.5\cdot b$ respectively, for which we let $b$ vary between 0 and 0.5 in steps of 0.05.
+For each value of $a$ and each value of $b$ we generate 100 datasets, and calculate the mean AUROC across these datasets.
 Table 2: Real-life datasets used in the experiment. $m$: number of attributes.
 Dataset Source Records Anomalies $m$
-ai4i2020 [15] 10 000 Simulated machine operation records 339 Five different failure modes (not distinguished by the label) 6
-diabetes-risk [8] 520 Patients with diabetes symptoms 320 Patients actually diagnosed with diabetes 14
+ai4i2020 [15] 10 000 Simulated machine operation records 339 Five different failure modes 6
+diabetes-risk [8] 520 Patients with diabetes symptoms 320 Actual diabetes patients 14
 fertility [6] 100 Sperm samples 12 ‘Altered’ samples 8
-heart-failure [1] 299 Heart patients 96 Patients who died in the follow-up period 11
+heart-failure [1] 299 Heart patients 96 Patients who died 11
 phishing-websites [24] 11 055 Websites 4898 Phishing websites 30
@@
 qualitative-bankruptcy [9] 250 Companies 107 Companies that went bankrupt 6
-south-german-credit [6] 1000 Credits provided by a bank 300 Bad credits 20
+south-german-credit [7] 1000 Bank credits 300 Bad credits 20
 thoraric-surgery [28] 470 Primary lung cancer patients who underwent major lung resections 70 Patients who died within one year 16
@@
 wisconsin [26] 683 Fine-needle aspirates of breast tumors 239 Malignant tumors 9
-wpbc [20] 138 Patients who have undergone surgery for an invasive malignant breast cancer tumor with no evidence of distant metastases 28 Patients who experienced recurrence within two years 32
-Table 3: Mean cross-validation AUROC for the absolute, ramp and signed distance variants. Bold: highest value (before rounding) for NND and ALP, respectively.
-Dataset NND ALP
-Absolute Ramp Signed Absolute Ramp
-ai4i2020 0.823 0.922 0.724 0.877 0.924
-diabetes-risk 0.971 0.923 0.716 0.895 0.926
-fertility 0.602 0.653 0.540 0.581 0.636
-heart-failure 0.715 0.769 0.735 0.734 0.766
-phishing-websites 0.901 0.927 0.804 0.927 0.936
-post-operative 0.476 0.504 0.557 0.459 0.484
-qualitative-bankruptcy 1.000 1.000 0.998 1.000 1.000
-south-german-credit 0.648 0.718 0.683 0.648 0.714
-thoraric-surgery 0.597 0.624 0.583 0.634 0.621
-wdbc 0.950 0.976 0.969 0.957 0.981
-wisconsin 0.995 0.994 0.995 0.872 0.995
-wpbc 0.570 0.625 0.633 0.537 0.654
-5 Experiments on real-life datasets
-We will now evaluate our proposal on a number of real-life datasets from the UCI repository [3] that can be approached as directional anomaly detection problems.
+wpbc [20] 138 Patients with surgery for invasive malignant breast cancer 28 Patients who experienced recurrence within two years 32
+5.3 Real-life datasets
+We have selected 10 real-life datasets from the UCI repository [4] that can be approached as monotonic anomaly detection problems.
 The main properties of these datasets are listed in Table 2.
-The attributes of these datasets express machine operating conditions (ai4i2020), medical symptoms, comorbidities and lifestyle factors (diabetes-risk, fertility, heart-failure, post-operative and thoraric-surgery), tumor characteristics (wdbc, wisconsin and wpbc), and risk indicators provided by experts (phishing-websites, qualitative-bankrupty and south-german-credit).
-ai4i2020 and post-operative contain a mixture of directional and adirectional attributes, the other datasets only have directional attributes.
-As in the previous section, we use AUROC to evaluate anomaly detection performance.
-We perform 5-fold cross-validation on the normal records from each dataset, creating at each iteration a test set by combining one fifth of the normal records with all anomalous records.
-For NND, we use linear weights and fix $k=8$.
-We rescale all attributes by subtracting the midhinge and dividing by the semi-interquartile range of the normal training values, such that the interquartile range becomes $[-1,1]$ in the training data.
-Table 3 lists the mean cross-validation AUROC obtained by the various variants for each dataset.
-The first thing to note is that for NND, unlike what we saw for synthetic data in the previous section, ramp distance generally performs (much) better than signed distance.
-The only datasets for which signed distance performs better are post-operative and wpbc, for which the AUROC scores are very low anyway.
-When we perform a one-side Wilcoxon signed-rank test, we obtain $p=0.021$.
-We can conclude from this that real-life anomaly detection problems are different from our synthetic datasets in one important aspect: an unexpectedly low score on one risk factor does not compensate for high scores on other risk factors.
-However, these experiments do confirm the usefulness of directional anomaly detection, because ramp distance performs about as well or better than absolute distance.
-For NND, the signed-rank test gives $p=0.011$, or $p=0.023$ if we apply Holm-Bonferroni correction [7] for multiple testing, while for ALP it gives $p=0.0029$.
-There is one notable exception: diabetes-risk for NND.
+The attributes of these datasets express machine operating conditions (ai4i2020), medical symptoms, comorbidities and lifestyle factors (diabetes-risk, fertility, heart-failure, post-operative and thoraric-surgery), tumor characteristics (wdbc, wisconsin and wpbc), and risk indicators provided by experts (phishing-websites, qualitative-bankruptcy and south-german-credit).
+ai4i2020 and post-operative contain a mixture of monotonic and ordinary attributes, the other datasets only have monotonic attributes.
+On these datasets, we perform 5-fold cross-validation on the normal records, creating at each iteration a test set by combining one fifth of the normal records with all anomalous records.
+5.4 Hyperparameters
+The anomaly detection algorithms from Section 3 have a number of hyperparameters which can in principle be tuned when there are anomalies available for validation purposes [13].
+In general, however, this is not the case in the semi-supervised setting, so we will use sensible default values.
+For NND and LOF, 1-distance (Boscovich distance) generally outperforms 2-distance (Euclidean distance) [12, 11], and for ALP, 1-distance is the standard choice [12].
+For consistency, we also choose 1-distance for CD.
+For both weighted NND and ALP, the standard choice is to use linearly descending weights [12, 11].
+As for the number of neighbours, $k=8$ has been established as the optimal default value for weighted NND [11], $k=2.5\log n$ for LOF [12] and $k=5.5\log n$ and $l=6\log n$ for ALP [12], where $n$ is the number of training records.
+For CD, we will use the marginal mean (centroid) as the centre value.
+Finally, while not typically seen as a hyperparameter, all distance-based algorithms also depend on the relative scale of the attributes.
+We normalise the real-life datasets by dividing each attribute by the semi-interquartile range of the normal training values, such that the interquartile range becomes $[-1,1]$ in the training data.
+Figure 1: Mean AUROC obtained with synthetically generated datasets with Gaussian and Bernoulli attributes.
+$a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
+6 Results
+6.1 Synthetic datasets
+The results for the synthetic datasets are displayed in Figure 1.
+With Gaussian attributes, the results are substantially the same for the different anomaly detection algorithms.
+Ramp distance clearly performs better than absolute distance, in line with our monotonic hypothesis.
+For CD, signed distance performs even better than ramp distance, but the difference is small.
+With Bernoulli attributes, ramp distance also consistency performs well, but there is more variation for absolute distance.
+For ALP and NND, the performance deficit with respect to ramp distance is much smaller than with Gaussian attributes, and for CD all three distance measures give practically identical results.
+Finally, for LOF, the results with absolute distance are particularly poor.
+Taken together, these results suggest that when we know for a given attribute that anomalies have a higher mean value than normal records, we should indeed use ramp distance rather than absolute distance.
+On these idealised datasets, CD performs as well as the nearest neighbour algorithms, and signed distance appears to produce even better results than ramp distance.
+However, we will see in the next subsection that this does not carry over to the more complicated structure of real-life datasets.
+Table 3: Mean cross-validation AUROC for the absolute (A), ramp (R) and signed (S) distance variants. Bold: highest value (before rounding) for each algorithm.
+Dataset CD NND LOF ALP
+A R S A R A R A R
+ai4i2020 0.807 0.868 0.853 0.823 0.922 0.867 0.909 0.877 0.924
+diabetes-risk 0.886 0.778 0.715 0.971 0.923 0.745 0.920 0.895 0.926
+fertility 0.388 0.478 0.540 0.602 0.653 0.534 0.605 0.581 0.636
+heart-failure 0.714 0.773 0.735 0.715 0.769 0.736 0.715 0.734 0.766
+phishing-websites 0.665 0.767 0.804 0.901 0.927 0.851 0.902 0.927 0.936
+post-operative 0.481 0.522 0.560 0.476 0.504 0.522 0.527 0.459 0.484
+qualitative-bankruptcy 1.000 1.000 0.998 1.000 1.000 1.000 1.000 1.000 1.000
+south-german-credit 0.613 0.711 0.683 0.648 0.718 0.638 0.675 0.648 0.714
+thoraric-surgery 0.574 0.629 0.583 0.597 0.624 0.607 0.542 0.634 0.621
+wdbc 0.953 0.969 0.969 0.950 0.976 0.950 0.954 0.957 0.981
+wisconsin 0.995 0.995 0.995 0.995 0.994 0.692 0.990 0.872 0.995
+wpbc 0.529 0.624 0.633 0.570 0.625 0.555 0.582 0.537 0.654
+6.2 Real-life datasets
+Table 3 lists the mean cross-validation AUROC obtained by the various variants for each real-life dataset.
+CD does not perform quite as well as the other algorithms, which is not surprising given the more complicated structure of these datasets.
+However, it is remarkable that it performs as well as it does.
+Overall, NND and ALP achieve the best performance and consistently outperform LOF.
+The main question that we want to answer is whether monotonic anomaly detection outperforms ordinary anomaly detection on monotonic datasets.
+Note first that contrary to what we saw for synthetic datasets, signed distance no longer outperforms ramp distance for CD.
+Moreover, it is consistently beaten by the combination of NND with ramp distance.
+This leaves ramp distance as the clearly preferred monotonic variant.
+In line with our monotonic hypothesis, we find that ramp distance significantly outperforms absolute distance: the one-sided Wilcoxon signed-rank test gives $p=0.030$ (CD), $p=0.011$ (NND), $p=0.025$ (LOF) and $p=0.0029$ (ALP).
+In those cases where absolute distance performs better than ramp distance, the difference is very small or the AUROC scores are altogether very low.
+There is one notable exception: the diabetes-risk dataset for CD and NND.
 A potential explanation for this is the fact that the attributes of this dataset encode diabetes symptoms, and that moreover, the normal records do not represent average healthy people, but non-diabetes patients who nevertheless display symptoms of diabetes.
-Consequently, this may undermine the directionality of this dataset — diabetes patients may not necessarily have more symptoms than this particular group of other patients, but rather different combinations of symptoms, which would make absolute distance a better fit than ramp distance.
+Consequently, this may undermine the monotonicity of this dataset — diabetes patients may not necessarily have more symptoms than this particular group of other patients, but rather different combinations of symptoms, which would make absolute distance a better fit than ramp distance.
 Indeed, when we look at the mean attribute values of normal and anomalous records in this dataset, we find that for five of the fourteen attributes, the anomalous records (diabetes patients) have lower or only slightly higher mean values than the normal records.
-If we rerun the experiment while treating these five attributes as non-directional, the AUROC obtained by ramp distance matches that of absolute distance.
+If we rerun the experiment while treating these five attributes as non-monotonic, the AUROC obtained by ramp distance matches that of absolute distance.
 7 Conclusion
-In this paper, we have introduced directional anomaly detection, a new problem setting wherein only high (or only low) values of certain attributes are indicative of anomality, and which can therefore be viewed as risk factors.
-We have proposed two different ways of adapting distance-based anomaly detection algorithms to make use of this knowledge: ramp distance and signed distance.
-The difference between these two variants corresponds to a choice in how a practitioner wants to interpret an unexpectedly low value on a given risk factor.
+In this paper, we have introduced monotonic anomaly detection, a new problem setting that incorporates the domain knowledge that only high (or only low) values of certain attributes are indicative of anomality (and which can therefore be viewed as risk factors).
+We have shown that distance-based anomaly detectors can take this monotonicity into account by using an asymmetric distance function.
+For this purpose, we have proposed two variants of the commonly used family of Minkowski $p$-distances: ramp distance and signed distance.
+The difference between these two variants corresponds to a choice in how a practitioner wants to interpret a relatively low value of a given attribute.
 With ramp distance, such low values are simply discounted, whereas with signed distance, they contribute negative evidence against anomality.
-Both variants contrast with absolute distance, the non-directional baseline wherein such low values contribute positive evidence for anomality.
-In an experiment with synthetically generated Gaussian and Bernoulli data, we found that signed distance is able to achieve slightly higher anomaly detection than ramp distance.
-However, in a subsequent experiment with real-life datasets, ramp distance performed substantially better.
-We conclude from this that in typical real-life use cases, unexpectedly low values on risk factors should not compensate for high values on other risk factors.
-We also found that ramp distance generally performed as good or better than absolute distance.
-Therefore, we recommend the use of ramp distance when a practitioner knows that some of the attributes in their dataset are really risk factors.
-Finally, we will end by briefly considering what it means if, for a given dataset, ramp distance should nevertheless result in worse anomaly detection performance than absolute distance.
-One possibility is that the directional hypothesis is false for one or more of the attributes, that there is a good reason after all why (some) anomalies should have lower values than normal.
-This is in fact what we saw with one of the real-life datasets that we evaluated, the only dataset where absolute distance clearly performed better than ramp distance for NND.
-The other possibility is that there is no clear explanation, no plausible causal relation linking a lower attribute value to a higher risk of abnormality.
-In that case using ramp distance may still be preferable, because it better matches the domain knowledge of the practitioner, resulting in a more interpretable prediction model.
-Acknowledgements
+Both variants contrast with absolute distance, the non-monotonic baseline wherein such low values contribute positive evidence for anomality.
+However, because signed distance lets low values of some attributes compensate for high values of other attributes, the precise distribution of the attribute values becomes irrelevant, and anomality is purely determined on the basis of the total attribute value sum.
+This makes signed distance fundamentally incompatible with nearest neighbour algorithms, which are based on the premise that the normal data class is characterised by a specific distribution of attribute values.
+This contrast is reflected in our results on synthetic and real-life data.
+On the idealised synthetic datasets, the simple CD algorithm performed about as well as the nearest neighbour algorithms, and signed distance achieved slightly higher performance than ramp distance.
+However, on the more complex real-life datasets, the nearest neighbour algorithms with ramp distance outperformed CD with signed distance.
+Therefore, ramp distance is the preferred monotonic variant.
+When comparing ramp distance with absolute distance, which corresponds to ordinary anomaly detection, we found that on both synthetic and real-life datasets, ramp distance performs significantly better.
+We conclude that practitioners faced with a monotonic anomaly detection problem — where they know that only higher values of certain attributes are indicative of anomality — should use a nearest neighbour algorithm that uses ramp distance for monotonic attributes, for two reasons.
+Firstly, because the resulting model aligns better with the domain knowledge of the practitioner.
+And secondly, because it is likely to lead to higher anomaly detection performance.
+Finally, we note that there are other types of anomaly detection algorithms, like the Support Vector Machine (SVM) and the Isolation Forest (IF), which should also be adapted for monotonic anomaly detection.
+{credits}
+7.0.1 Data and code.
+Datasets and the code to reproduce our experiments are available at https://liacs.leidenuniv.nl/~lenzou/code/lenz-2025-monotonic.tar.gz.
+7.0.2 Acknowledgements
 This publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO).
+Figure 2: Mean AUROC for NND as a function of $k$, across real-life datasets.
+Appendix 0.A The number of neighbours $k$
+In semi-supervised anomaly detection, we generally do not have access to anomalies for the purpose of hyperparameter tuning.
+Therefore, we have chosen default values in our main experiment.
+Nevertheless, we may ask how performance on our test sets varies as a function of different hyperparameter values.
+For reasons of space, we limit ourselves here to the number of neighbours $k$ of NND.
+Fig.
+2 displays the mean AUROC of ramp and absolute distance as a function of $k$, for the real-life datasets.
+We observe that the default value of $k=8$ lies close to the maximum for both distance variants, providing validation that this is a good default value.